Find the square roots of the following numbers by the Prime Factorisation Method.   vi. 9604  

Grade:12th pass

1 Answers

Harshit Singh
2 years ago
Dear Student

9604=(2×7×7)×(2×7×7) ⇒9604=(2×7×7)2⇒√9604=2×7×7=98
Thanks